Find the inverse of the following matrix, using Gauss Jordan method.
1 2 3 4
2 3 4 1
3 4 1 2
4 1 2 3
1
Expert's answer
2020-04-21T15:33:58-0400

"A=\begin{bmatrix}
 1 & 2&3&4 \\
 2&3&4&1\\
3&4&1&2\\
4&1&2&3
\end{bmatrix}"

"[A|E]=\begin{bmatrix}
 1 & 2&3&4 |1&0&0&0\\
 2&3&4&1|0&1&0&0\\
3&4&1&2|0&0&1&0\\
4&1&2&3|0&0&0&1
\end{bmatrix}"

IIr+Ir(-2)

IIIr+Ir(-3)

IIIIr+Ir(-4)

"\begin{bmatrix}
 1 & 2&3&4 |1&0&0&0\\
 0&-1&-2&-7|-2&1&0&0\\
0&-2&-8&-10|-3&0&1&0\\
0&-7&-10&-13|-4&0&0&1
\end{bmatrix}"

IIr(-1)

IIIr+IIr(-2)

IIIIr+IIr(-7)

"\begin{bmatrix}
 1 & 2&3&4 |1&0&0&0\\
 0&1&2&7|2&-1&0&0\\
0&0&-4&4|1&-2&1&0\\
0&0&4&36|10&-7&0&1
\end{bmatrix}"

IIIr"\cdot(-\frac{1}{4})"

IIIIr+IIIr

"\begin{bmatrix}
 1 & 2&3&4 |1&0&0&0\\
 0&1&2&7|2&-1&0&0\\
0&0&1&-1|-0.25&0.5&-0.25&0\\
0&0&0&40|11&-9&1&1
\end{bmatrix}"

IIIIr"\cdot\frac{1}{40}"

IIIr+IIIIr

IIr+IIIr(-2)+IIIIr(-7)

Ir+IIr(-2)+IIIr(-3)+IIIIr(-4)

"\begin{bmatrix}
 1 & 0&0&0 |\\&&&&-0.225&0.025&0.025&0.275\\
 0&1&0&0|\\&&&&0.025&0.025&0.275&-0.225\\
0&0&1&0|\\&&&&0.025&0.275&-0.225&0.025\\
0&0&0&1|\\&&&&0.275&-0.225&0.025&0.025
\end{bmatrix}"

"A^{-1}==\begin{bmatrix}
 -0.225&0.025&0.025&0.275\\
0.025&0.025&0.275&-0.225\\
0.025&0.275&-0.225&0.025\\
0.275&-0.225&0.025&0.025
\end{bmatrix}"